OpenStudy (anonymous):

What are the solutions of the quadratic equation? 4x^2+34x+60=0

OpenStudy (anonymous):

polynomial of form: ax^2 + bx + c = 0 \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] a = 4 b = 34 c = 60

OpenStudy (anonymous):

let me know if this needs clarifications ^_^

OpenStudy (anonymous):

Thank you! I got it! :) @Euler271

OpenStudy (anonymous):

^_^